Solano Company has sales of $740,000, cost of goods sold of $490,000, other operating expenses of $46,000, average invested assets of $2,200,000, and a hurdle rate of 10 percent.

Required:
1. Determine Solano’s return on investment (ROI), investment turnover, profit margin, and residual income.
2. Several possible changes that Solano could face in the upcoming year follow. Determine each scenario’s impact on Solano’s ROI and residual income. (Note: Treat each scenario independently.)
a. Company sales and cost of goods sold increase by 30 percent.
b. Operating expenses decrease by $10,000.
c. Operating expenses increase by 10 percent.
d. Average invested assets increase by $420,000.
e. Solano changes its hurdle rate to 16 percent.
Answer #1

1.

sale

740,000

Less: Cost of goods sold

490,000

gross profit

250,000

operating expense

46000

net income

204,000

ROI = net income /average operating assets = 204000/2,200,000 = 9.27%

Investment turnover = sales /Average operating assets = 740,000/2,200,000 = 0.34 times

Profit margin = net income /sales = 204000/7400000 = 27.6%

Residual income = net income - ( operating assets * hurdle rate) = 204,000 - (2200,000*10%) = -16000

2. a.

Sales

962000

cost of goods sold

637000

Gross profit

325000

operating expense

46000

Net income

279000

ROI   = 279000/2,20,000 =12.68%

Residual income = 279,000 - (2200,000*10%) = 59,000

b.

Cost of goods sold

740,000

Cost of goods sold

490,000

gross profit

250,000

operating expense

36,000

net income

214,000

ROI = 214000/2,200,000 = 9.73%

Residual income = 214000 - 220000 = -6000

c.

gross profit

250,000

operating expense

41400

net income

208,600

ROI = 208600/2,200,000 = 9.48%
Residual income =208600 - 220,000 = -11,400

d. ROI = 204000/2,620,000*100 = 7.79%   

Residual loss =   204,000 - (2,620,000*10%) = -58,000

e. Residual loss = 204,000 - (2,200,000*16%) = -148,000